Dear Richard dear all,

please find below a short summary of the activities done today on the mountain related to the AO system. Italian people working at LBTO were:

Arcetri: E. Pinna, A. Puglisi, L. Fini, M. Sozzi, P. Stefanini, S. Esposito, C. Arcidiacono, A. Riccardi, M. Xompero, R. Briguglio
MG: R. Biasi

Total # of person: 11

W unit

1) The W unit spare parts and tool boxes arrived to the telescope.
2) The telecentric lens was installed in the W unit.

AO control SW

1) M. Xompero continues config adsec sw. Fixed configuration problems and reported in TODO list.
2) FL communication test fails.

LBT672a

1) The test of the capacitive sensors using an external capacitance is in progress.
2) The process to unpack TS3 was started today. It was still ongoing at 3:00PM.

Logistic

1) Due to the deep snow on the mountain road most the Italian group reached the telescope at 12:00. The group then left the mountain at 15:00 to come down before dark.

2) G. Brusa & R. Souza joined the Arcetri team to unpack and install the thin shell TS3.

Next update tomorrow.

All the best

Simone

-- ArmandoRiccardi - 02 Apr 2010
Topic revision: r1 - 02 Apr 2010, ArmandoRiccardi
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback